OpenStudy (anonymous):

Bob has a tree house at a height of 20 ft. When he's feeling adventurous he uses a rope to climb the tree house instead of the ladder. If the weight of the rope is 32 lb, then how much work must Bob do to pull the rope up once he has entered the house?

OpenStudy (anonymous):

Work = force x distance

OpenStudy (anonymous):

I can't get the right answer. what is it?

OpenStudy (anonymous):

We will need to integrate from 0 to 20 ...

OpenStudy (anonymous):

I got 865 but that is not one of my answer choices

OpenStudy (asnaseer):

@claire7 - how did you get 865?

OpenStudy (anonymous):

w=fd..what is the right answer?

OpenStudy (asnaseer):

could you please explain how you got 865?

OpenStudy (anonymous):

by converting 32lb to N and multiplying what the conversion was which was 142 anf thenI did w=142(20) I am so confused

OpenStudy (asnaseer):

you have left out the acceleration due to gravity

OpenStudy (jamesj):

A bit messy but this is what you need to do:

OpenStudy (asnaseer):

F=ma

OpenStudy (asnaseer):

you also need to convert feet to meters to get the correct units

OpenStudy (turingtest):

we need the right formula for the weight, not just use 32, because that is a function of how much rope is hanging out of the tree house James' method covers that

OpenStudy (jamesj):

To explain my attachment: You have a small piece of rope, dx long at a vertical distance x from the tree house. How much work is required to move that piece of rope to the tree house? Well, it's the weight of that piece of rope times the distance covered. How much mass does that piece dx have? Well there 32 lbs in 20 ft, so for a distance dx of rope, it has a mass of \[ \frac{32}{20} dx \] Hence the weight of that piece of rope is \[ \frac{32}{20} g \ dx \] and therefore the work required to move that piece of rope to the tree house is \[ \frac{32}{20} g x \ dx \]

OpenStudy (jamesj):

Now, add up all of those pieces of work over the length of the rope, and the total work required to get the rope in the tree house is \[ W = \int_0^{20} \frac{32}{20} gx \ dx \]

OpenStudy (asnaseer):

@JamesJ - and then convert to the desired units?

OpenStudy (jamesj):

Yes, this is in American units. I'm guessing we can and probably should leave it in these units as this was what was given to us. I'm an SI man myself, but I'd roll with these units for now.

OpenStudy (asnaseer):

we use SI here as standard :-)

OpenStudy (jamesj):

Everywhere else in the world my friend. I wish the Americans would get with the program. Anyway, @claire, make sense?

OpenStudy (anonymous):

kind of you are a genius and I am not would the answer be somehwere around 200?

OpenStudy (jamesj):

Not quite. Leaving g as g in the integral, what do you get when you integrate?

OpenStudy (jamesj):

i.e., what is the anti-derivative of (32/20)gx ?

OpenStudy (anonymous):

15.6?

OpenStudy (jamesj):

\[ \int_0^{20} \frac{32g}{20} x \ dx = \frac{32g}{20} \int_0^{20} x \ dx \] What's the value of \[ \int_0^{20} x \ dx \]

OpenStudy (anonymous):

that does not look english to me I have not learned any of that yet

OpenStudy (turingtest):

no calculus allowed?

OpenStudy (jamesj):

Ok, well you'll get the same result if you think of moving up the center of mass (CM) of the rope. At what distance is the CM from the tree house?

OpenStudy (jamesj):

Where is the center of mass, CM, in the rope?

OpenStudy (jamesj):

The CM is exactly in the middle of the rope. Given that, how far does the CM have to move vertically to get into the tree house?

OpenStudy (asnaseer):